OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = –35 ?

OpenStudy (anonymous):

a17 = 13 + 16d = -35 16d = -48 d = -3

OpenStudy (anonymous):

Therefore we can find a35 = 13 + 34(-3) = 115